OpenStudy (anonymous):

I need to get x by itstelf:y= x^3 +3sinx+2cosx

ganeshie8 (ganeshie8):

who gave u this problem ?

OpenStudy (dumbcow):

not possible algebraically

OpenStudy (anonymous):

It is a calc 2 problem. My calc is not that strong that is hwy i was a bit confused too. If you guys know anything about inverse functions comment back.

ganeshie8 (ganeshie8):

is that the entire problem or you ended up with this while solving some other problem ?

OpenStudy (anonymous):

well in addition to that stuff. a=2 and i need to find (f^-1)'(a)

OpenStudy (dumbcow):

you can use this: http://en.wikipedia.org/wiki/Inverse_function#Inverses_and_derivatives \[\frac{dy}{dx} = 3x^2 +3\cos x - 2 \sin x\] thus \[\frac{dx}{dy} = (f^{-1})' (y) = \frac{1}{3y^2 +3\cos y -2\sin y}\]